OpenStudy (anonymous):

Find the numerical value of the rational expression when (a) x=4 (b) x=-2 x^2-4 / 3x+4

OpenStudy (anonymous):

a= (4)^2-4 16-4=12 / 3(4)+4 12+4=16 12/16 = 3/4 a=3/4

OpenStudy (anonymous):

b= (-2)^2-4 = 0 3(-2)+4 -6+4 = -2 b=0

OpenStudy (anonymous):

thank you!!!!